(a) the superficial mass velocity of air through a dehumidifier is found to be 300 lb/h.ft2 . calculate its equivalent value in kg/s.m2. (b) prove that 750.06 mm of Hg is equal to 1 bar (c) 150 g of carbon di oxide (Co2) is available in a tank of volume 100 cm3 at a temperature of 100 F . calculate the pressure of the gas in mm of Hg.
1
Expert's answer
2020-05-06T13:59:00-0400

(a) the equivalent mass velocity is 300 lb/(h*ft2) * 0.4536 kg/lb / 3600 s/h / (0.3048 m/ft)2 =

0.4069 kg/(s*m2)

(b) By definitions:

760 mmHg = 1 atm = 101325 Pa

1 bar = 100000 Pa = 1/1.01325 atm = 760/1.01325 mmHg = 750.06 mmHg

(c) Suppose ideal gas, when:

p*V = m/M*R*T

m = 150 g,

M = 44 g/mole,

R = 8.314 J/(K*mole),

T = 5/9*(100 F - 32) + 273.15 = 310.93 K,

V = 100 cm3 * 10-6 m3/cm3 = 10-4 m3

p = 150/44*8.314*310.93/10-4 = 88127455 Pa = 88127455 Pa / 101325 Pa/atm * 760 mmHg/atm = 661010 mmHg

Answer: pressure is 661010 mmHg